You were hired as a consultant to Quigley Company, whose target capital structure is 35% debt, 10% preferred, and 55% common equity. The interest rate on new debt is 6.50%, the yield on the preferred is 6.00%, the cost of retained earnings is 9.25%, and the tax rate is 40%. The firm will not be issuing any new stock. What is Quigley's WACC? Round final answer to two decimal places. Do not round your intermediate calculations.

Group of answer choices

7.05%

6.56%

5.57%

7.19%

8.25%
Answer #1

After-tax cost of debt=yield to maturity*(1-tax rate)

=6.5*(1-0.4)=3.9%

WACC=Respective cost*Respective weight

=(0.35*3.9)+(0.1*6)+(0.55*9.25)

=7.05%(Approx).